OpenStudy (anonymous):

MEDAL!! sin(theta/2)=1/2. Solve for theta and give a general formula for all the solutions

OpenStudy (anonymous):

I know the answers are pi/3 and 5pi/3

OpenStudy (anonymous):

but I don't know the general formula

OpenStudy (ksaimouli):

sin(pi/3)=1/2 and sin(5pi/3)=1/2 so theta= pi/3+k*pi

OpenStudy (ksaimouli):

past question that is not the general formula instead that is the basic structure u need to plug in theta like ^ one

OpenStudy (anonymous):

wait, what is the general formula for both? for the last question I just wrote tan(x)=k(pi)+a

OpenStudy (ksaimouli):

noooo as I said write like (^) one

OpenStudy (anonymous):

I'm so confused lol what was the general formula for the last question? the equation was tan theta=-1

OpenStudy (anonymous):

@ksaimouli

OpenStudy (ksaimouli):

okay what r th solution for the (^) one

OpenStudy (anonymous):

for tan theta=-1 the solutions were 3pi/4 and 7pi/4

OpenStudy (ksaimouli):

okay so theta= pi/4+kpi

OpenStudy (anonymous):

so theta=pi/4+kpi is the general formula?

OpenStudy (anonymous):

?? and for this question is he formula the one you wrote above?

OpenStudy (ksaimouli):

yes

OpenStudy (ksaimouli):

what math is this?

OpenStudy (anonymous):

pre-calc/trig

OpenStudy (ksaimouli):

when I took it nobody asked us to write the general formula